OpenStudy (abbycross167):

Can someone please try to explain to me how to answer this question? Please I will fan and give a medal!!! Question~ Solve the systems using the substitution method. y = (2/3)x + 1 y + x = 6

OpenStudy (abbycross167):

@Hero @pooja195 @dan815 @Data_LG2 @Luigi0210 could one of y'all please try to help me? I really don't know what to do with the (2/3)

OpenStudy (neonumbrella5115):

When you use substitution you substitute one value, say y, into another equation. In your problem we will substitute the equation equal to y into the second equation.

OpenStudy (abbycross167):

I think I did that step but after that, I don't know what to do. This is as far as I got~ . y = (2/3)x + 1 y + x = 6 Solve one equation for ether variable ~ Y+x=6 = (2/3)x+1+x=6